# Differential Contrastive Training for Gaze Estimation

### v2

@@ Abstract
The complex application scenarios have raised critical requirements for precise and generalizable gaze estimation methods. Recently, the pre-trained CLIP has achieved remarkable performance on various vision tasks, but its potentials have not been fully exploited in gaze estimation. In this paper, we propose a novel Differential Contrastive Training strategy, which boosts gaze estimation performance with the help of the CLIP. Accordingly, a Differential Contrastive Gaze Estimation network (DCGaze) composed of a Visual Appearance-aware branch and a Semantic Differential-aware branch is introduced. The Visual Appearance-aware branch is essentially a primary gaze estimation network and it incorporates an Adaptive Feature-refinement Unit (AFU) and a Double-head Gaze Regressor (DGR), which both help the primary network to extract informative and gaze-related appearance features. Moreover, the Semantic Difference-aware branch is designed on the basis of the CLIP’s text encoder to reveal the semantic difference of gazes. This branch could further empower the Visual Appearance-aware branch with the capability of characterizing the gaze-related semantic information. Extensive experimental results on four challenging datasets over within and cross-domain tasks demonstrate the effectiveness of our DCGaze. Code will be available upon acceptance.
1 Introduction
Gaze estimation aims to predict a 2D gaze position or a 3D gaze direction of a specified user given its facial or eye image. As an important task in the field of computer vision, it has been extensively utilized in various people-related researches, such as virtual reality [4, 23], autonomous driving [28, 11], etc. In those real scenarios, the users’ identities and their surrounding environments are changeable and complicated, which puts a heavy responsibility on the accuracy and generalization of gaze estimation models. In other words, the estimation models that are optimized via training users should adapt to disjointed new subjects accurately.
@@ 1 Introduction
Figure 1: (a) The gaze errors of ResNet18, ResNet50 and CLIP on MPIIFaceGaze dataset. We evaulate ResNet18 and ResNet50 on the cross-domain task $\mathcal{D}_{E}\rightarrow\mathcal{D}_{M}$, which indicates training on ETH-XGaze and testing on MPIIFaceGaze. (b) Comparing with generating a unique sentence for each image, it is more easier to describe the gaze differences of a pair of images.
@@
Most recently, the pre-trained Visual-Language models represented by CLIP [31] have demonstrated impressive performance on various perception tasks, like image classification [12], semantic segmentation [55], depth estimation [47, 46] and others [17]. CLIP trains an image encoder and a text encoder jointly by a contrastive loss in the embedding space between both modalities, which endows it with powerful representation capabilities by characterizing both visual and semantic information of a target object. Naturally, we are spontaneously curious about the issue: can CLIP understand human gaze? To explore this concern, we make a simple early attempt. Inspired by [47], we define four gaze bins that align with the gaze directions of $[0,\frac{\pi}{2}]$, $[0,-\frac{\pi}{2}]$, $[\frac{\pi}{2},0]$ and $[-\frac{\pi}{2},0]$, respectively. And their semantic descriptions are designated as a prescribed prompt ‘A photo of a face gazing {gaze direction}.’ where gaze direction includes [‘up’, ‘down’, ‘left’, ‘right’]. Then the gaze direction of an arbitrary facial image could be obtained via linearly combining the multi-bin gaze values according to the language-image similarities between its visual embedding and semantic embeddings of all the gaze bins (details refer to supplementary materials). Surprisingly, even without fine-tuning on any exclusive gaze estimation datasets, the CLIP achieves promising performance comparable to ResNet18 and ResNet50 that trained with cross-domain samples (Figure 1). From the results, we thus draw a rough conclusion that the CLIP could perceive human gazes subtly, but its potentials have not been fully exploited.
@@
Recently, several studies [45, 38] have attempted to take advantage of CLIP to predict gaze directions. The notable CLIP-Gaze [45] aims to purify gaze-relevant features by pushing away the extracted gaze features from gaze-irrelevant semantic embeddings obtained via the text encoder of CLIP. Essentially, it employs the text encoder to generate gaze distractors rather than describe the gaze itself, which remains unable to solve the above issues.
@@
To address the aforementioned challenges, in this paper, we put forward an innovative and straightforward Differential Contrastive Training (DCTrain) strategy to fully evoke CLIP’s potentials in gaze estimation. Commonly, although it is impractical to describe the continuous gaze direction of each facial image in language, it is easier to semantically distinguish the gaze difference between a sample pair. For example, as shown in Figure 1, there are three facial images with gaze directions of $[0.54,-0.20]$, $[0.24,-0.43]$, $[0.08,0.42]$, respectively. Obviously, we could not generate a unique sentence for each image to explain their detailed gaze directions. While we could clearly draw the conclusions that ‘the people in image1 and image2 are looking in much similar directions’ and ‘the people in image1 and image3 are looking in much different directions.’ Motivated by the common observations, instead of identifying gaze direction of single facial image, our Differential Contrastive Training strategy is designed to leverage CLIP for unveiling gaze semantic differences of image-pairs. Innovatively, this strategy could simplify the target texts that the CLIP needs to understand by replacing its unfamiliar phrases, e.g., ‘$[0.54,-0.20]$’ with familiar ones, e.g., ‘similar’ or ‘different’ (Figure 1). Moreover, the continuous gaze labels are transformed into discrete similarity levels of a pair of gazes. Thus, the above two challenges are expected to be resolved under our Differential Contrastive Training strategy.
Accordingly, a Differential Contrastive Gaze Estimation network (DCGaze) is proposed, which consists of a Visual Appearance-aware branch and a Semantic Difference-aware branch.
@@
The Visual Appearance-aware branch is essentially a primary gaze estimation network, which extracts the appearance feature of each facial image and then regresses it into a final gaze direction. We employ the remarkable CNN-Transformer architecture as the basic structure of this branch. To ensure that the extracted appearance features are informative and gaze-related, we propose an Adaptive Feature-refinement Unit (AFU) and a Double-head Gaze Regressor (DGR). The former is based on the image encoder of the pre-trained CLIP, which aims to dynamically enhance the gaze-related contents contained in extracted appearance features via thoroughly exploring its relationships with the prior appearance features of CLIP. The Double-head Gaze Regressor is introduced to mitigate the overfitting issues caused by the conventional MLP-based regressor. It incorporates an extra regressor head with random masking policy to prevent itself from excessively focusing on several feature dimensions.
The Semantic Difference-aware branch is built upon the text encoder of CLIP, which could be plug and play to arbitrary gaze estimation models. Firstly, a series of differential gaze prompts are designed, which reveal the relationships of gaze directions between two facial images in language (Table  1). And then, the connections between the semantic embeddings of these prompts and the appearance features of image-pairs captured via the Visual Appearance-aware branch could be established to realize an ‘image-language’ contrastive learning. To be specific, the ‘image’ refers to the concatenated appearance features of selected image-pairs. The ‘language’ refers to the sentence that describes their similarity in terms of true gaze directions. Therefore, by semantically identifying the gaze difference of sample pairs with the help of the Differential Contrastive Training, we can fully leverage the semantic representation ability of CLIP to facilitate the primary gaze estimation model to characterize semantic gaze-related information.
@@
$\bullet$ We develop the potentials of CLIP in boosting gaze estimation performance with a novel Differential Contrastive Training strategy. Innovatively, we propose to establish the connections between languages and gaze differences of image -pairs, which could provide new ideas for analogous quantified/regression tasks.
@@
Figure 2: Framework of our proposed DCNet. It consists of two branch: Visual Appearance-aware branch and Semantic Difference-aware branch. In Visual Appearance-aware branch, the AFU is designed to adaptively enhance the gaze-related contents of primary appearance features with the guidance of the prior appearance features of CLIP’s image encoder. The DGR maps those enhanced features to final gazes via two regression heads.The Semantic Difference-aware branch selects several image-pairs from sample batch and gives each of them a textual prompt that describes their gaze differences. Then, the prompts are fed into CLIP’s text encoder to capture text embeddings, which are then aligned with enhanced appearance features obtained by our Visual Appearance-aware branch.
2 Related Work
@@
Gaze estimation. With the recent developments in deep learning, appearance-based gaze estimation methods have become the mainstream. Researchers explored various CNN-based architectures [50, 8, 6] to build the mapping between facial images and gaze directions. Zhang et al. [49] firstly proposed a CNN-based gaze estimation network and the well-known MPIIFaceGaze dataset. With the introduction of Transformer [36], the Vision Transformer-based structures have started to be applied to gaze estimation and achieved promising performance (e.g., GazeTR [7], oh et al.  [29], SUGE [39]). In the recent years, researchers dedicated themselves to inventing generalized gaze estimation methods, which would show robust performance on unseen users. Several methods leveraged a gaze redirection strategy to extend the datasets for generalized gaze estimation [16, 44]. Cheng et al. [10] introduced a method of purifying gaze features to improve the network’s generalization. Besides, some methods based on contrastive learning [41]and uncertainty learning [53, 39] also demonstrated remarkable generalizability. In this paper, we leverage CLIP to improve discriminability and generalization of the gaze estimation network.
@@
Pre-trained Vision-language Models. Recently, the CLIP [31] trained on large-scale image-text pairs have attracted increasing attentions. Because of its powerful visual and semantic representation capabilities, CLIP has been transferred to various vision tasks, such as objection detection [37], semantic segmentation [55], and others [25, 18, 42]. Especially, CLIP also shows surprising capacities on quantified vision tasks, e.g., depth estimation [47, 46], 3D hand poses estimation [17], etc. Moreover, researchers were attempting to take advantages of the pre-trained CLIP to predict gaze directions. Wang et al. [38] designed a linguistic description generator to produce text signals with coarse gaze directional cues. And then a cross-attention condenser was designed to finely recalibrate the visual and text representations, enhancing the learning quality of gaze features. Yin et al. [45] designed a feature separation loss by employing CLIP text encoder to generate gaze distractors from diverse language descriptions, which aims at purifying the gaze-relevant feature via pushing away it from gaze-irrelevant features. In this paper, we adapt CLIP to gaze estimation with an innovative collaborative enhancing strategy, in which the CLIP is regarded as an assistance to enhance the obtained gaze features.
3 Methodology
@@
To fully activate the potentials of CLIP to perceive gaze directions, we propose a novel Differential Contrastive Training (DCTrain) strategy, in which the CLIP is leveraged for unveiling gaze semantic differences of image-pairs. Accordingly, we propose a Differential Contrastive Gaze Estimation network (DCGaze) as shown in Figure 2 which consists of a Visual Appearance-aware branch and a Semantic Difference-aware branch. The former is essentially a primary gaze estimation network, while the latter is to encourage the Visual Appearance-aware branch to extract gaze-related features with the help of image-language contrastive learning.
3.1 Visual Appearance-aware Branch
@@
The Visual Appearance-aware branch is composed of a Basic Appearance Feature Extractor and two innovative modules, namely a Adaptive Feature-refinement Unit (AFU) and a Double-head Gaze Regressor (DGR). The basic feature extractor captures the appearance feature of each facial image. The two proposed modules are designed to help extract informative and gaze-related appearance features.
3.1.1 Basic Appearance Feature Extractor
@@
We employ the remarkable CNN-Transformer architecture [7] as the basic structure, where a ResNet is firstly adopted to acquire feature maps $f_{i}^{maps}\in\mathbb{R}^{W\times H\times C}$ of a given image $x_{i}\in X$ where $X=\{x_{1},x_{2},\dots,x_{n}\}$. Then those feature maps are reshaped into $W\times H$ patches $f_{i}^{p}\in\mathbb{R}^{\left(W\times H\right)\times C}$,which are treated as a series of $C$-dimensional visual tokens. After adding an extra learnable token $f_{i}^{token}\in\mathbb{R}^{1\times C}$, which is used to aggregate the features of all the patches, we feed them into a Transformer with a learnable position embedding $f_{i}^{pos}\in\mathbb{R}^{\left(1+W\times H\right)\times C}$. Overall, we get the final primary appearance feature $f_{i}^{pry}\in\mathbb{R}^{1\times C}$ as Eq. (1).
$f_{i}^{pry}=\text{Transformer}\left(\left[{f_{i}^{token}};{f_{i}^{p}}\right]+{f_{i}^{pos}}\right)\left[0,:\right],$ (1)
@@
where $[0,:]$ represents that the first row of the output feature maps is serving as the primary feature and [;][;][ ; ] denotes the concatenation operation.
3.1.2 Adaptive Feature-refinement Unit

### v3

@@ Abstract.
The complex application scenarios have raised critical requirements for precise and generalizable gaze estimation methods. Recently, the pre-trained CLIP has achieved remarkable performance on various vision tasks, but its potentials have not been fully exploited in gaze estimation. In this paper, we propose a novel Differential Contrastive Training strategy, which boosts gaze estimation performance with the help of the CLIP. Accordingly, a Differential Contrastive Gaze Estimation network (DCGaze) composed of a Visual Appearance-aware branch and a Semantic Differential-aware branch is introduced. The Visual Appearance-aware branch is essentially a primary gaze estimation network and it incorporates an Adaptive Feature-refinement Unit (AFU) and a Double-head Gaze Regressor (DGR), which both help the primary network to extract informative and gaze-related appearance features. Moreover, the Semantic Difference-aware branch is designed on the basis of the CLIP’s text encoder to reveal the semantic difference of gazes. This branch could further empower the Visual Appearance-aware branch with the capability of characterizing the gaze-related semantic information. Extensive experimental results on four challenging datasets over within and cross-domain tasks demonstrate the effectiveness of our DCGaze. The code is available at https://github.com/LinZhang-bjtu/DCGaze.
conference: ; ; copyright: none
@@
Figure 1. (a) The gaze errors of ResNet18, ResNet50 and CLIP on MPIIFaceGaze dataset. We evaulate ResNet18 and ResNet50 on the cross-domain task $\mathcal{D}_{E}\rightarrow\mathcal{D}_{M}$, which indicates training on ETH-XGaze and testing on MPIIFaceGaze. (b) Comparing with generating a unique sentence for each image, it is more easier to describe the gaze difference of a pair of images.
1. Introduction
Gaze estimation aims to predict a 2D gaze position or a 3D gaze direction of a specified user given its facial or eye image. As an important task in the field of computer vision, it has been extensively utilized in various people-related researches, such as virtual reality (Burova et al., 2020; Konrad et al., 2020), autonomous driving (Mole et al., 2021; Cheng et al., 2024), etc. In those real scenarios, the users’ identities and their surrounding environments are changeable and complicated, which puts a heavy responsibility on the accuracy and generalization of gaze estimation models. In other words, the estimation models that are optimized via training users should adapt to disjointed new subjects accurately.
@@ 1. Introduction
Most recently, the pre-trained Visual-Language models represented by CLIP (Radford et al., 2021) have demonstrated impressive performance on various perception tasks, like image classification (Cherti et al., 2023), semantic segmentation (Zhou et al., 2023), depth estimation (Zhang et al., 2022; Zeng et al., 2024) and others (Guo et al., 2023). CLIP trains an image encoder and a text encoder jointly by a contrastive loss in the embedding space between both modalities, which endows it with powerful representation capabilities by characterizing both visual and semantic information of a target object. Naturally, we are spontaneously curious about the issue: can CLIP understand human gaze? To explore this concern, we make a simple early attempt. Inspired by (Zhang et al., 2022), we define four gaze bins that align with the gaze directions of $[0,\frac{\pi}{2}]$, $[0,-\frac{\pi}{2}]$, $[\frac{\pi}{2},0]$ and $[-\frac{\pi}{2},0]$, respectively. And their semantic descriptions are designated as a prescribed prompt ‘A photo of a face gazing {gaze direction}.’ where gaze direction includes [‘up’, ‘down’, ‘left’, ‘right’]. Then the gaze direction of an arbitrary facial image could be obtained via linearly combining the multi-bin gaze values according to the language-image similarities between its visual embedding and semantic embeddings of all the gaze bins (details refer to supplementary materials). Surprisingly, even without fine-tuning on any exclusive gaze estimation datasets, the CLIP achieves promising performance comparable to ResNet18 and ResNet50 that trained with cross-domain samples (Figure 1). From the results, we thus draw a rough conclusion that the CLIP could perceive human gazes subtly, but its potentials have not been fully exploited.
@@
Recently, several studies (Yin et al., 2024; Wang et al., 2023) have attempted to take advantage of CLIP to predict gaze directions. The notable CLIP-Gaze (Yin et al., 2024) aims to purify gaze-relevant features by pushing away the extracted gaze features from gaze-irrelevant semantic embeddings obtained via the text encoder of CLIP. Essentially, it employs the text encoder to generate gaze distractors rather than describe the gaze itself, which remains unable to solve the above issues.
@@
To address the aforementioned challenges, in this paper, we put forward an innovative and straightforward Differential Contrastive Training (DCTrain) strategy to fully evoke CLIP’s potentials in gaze estimation. Commonly, although it is impractical to describe the continuous gaze direction of each facial image in language, it is easier to semantically distinguish the gaze difference between a sample pair. For example, as shown in Figure 1, there are three facial images with gaze directions of [0.54,0.20][0.54,-0.20][ 0.54 , - 0.20 ], [0.24,0.43][0.24,-0.43][ 0.24 , - 0.43 ], [0.08,0.42][0.08,0.42][ 0.08 , 0.42 ], respectively. Obviously, we could not generate a unique sentence for each image to explain their detailed gaze directions. While we could clearly draw the conclusions that ‘the people in image1 and image2 are looking in similar directions’ and ‘the people in image1 and image3 are looking in different directions.’ Motivated by the common observations, instead of identifying gaze direction of single facial image, our Differential Contrastive Training strategy is designed to leverage CLIP for unveiling gaze semantic differences of image pairs. Innovatively, this strategy could simplify the target texts that the CLIP needs to understand by replacing its unfamiliar phrases, e.g., ‘[0.54,0.20][0.54,-0.20][ 0.54 , - 0.20 ]’ with familiar ones, e.g., ‘similar’ or ‘different’ (Figure 1). Moreover, the continuous gaze labels are transformed into discrete similarity levels of a pair of gazes. Thus, the above two challenges are expected to be resolved under our Differential Contrastive Training strategy.
@@
Accordingly, a Differential Contrastive Gaze Estimation network (DCGaze) is proposed, which consists of a Visual Appearance-aware branch and a Semantic Difference-aware branch. The former is essentially a primary gaze estimation network, which extracts the appearance feature of each facial image and then regresses it into a final gaze direction. We employ the remarkable CNN-Transformer architecture as the basic structure of this branch. To ensure that the extracted appearance features are informative and gaze-related, we propose an Adaptive Feature-refinement Unit (AFU) and a Double-head Gaze Regressor (DGR). The former is based on the image encoder of the pre-trained CLIP, which aims to dynamically enhance the gaze-related contents contained in extracted appearance features via thoroughly exploring its relationships with the prior appearance features of CLIP. The Double-head Gaze Regressor is introduced to mitigate the overfitting issues caused by the conventional MLP-based regressor. It incorporates an extra regressor head with random masking policy to prevent itself from excessively focusing on several feature dimensions. The later branch is built upon the text encoder of CLIP, which could be plug and play to arbitrary gaze estimation models. Firstly, a series of differential gaze prompts are designed, which reveal the relationships of gaze directions between two facial images in language (Table  1). And then, the connections between the semantic embeddings of these prompts and the appearance features of image pairs captured via the Visual Appearance-aware branch could be established to realize an ‘image-language’ contrastive learning. To be specific, the ‘image’ refers to the concatenated appearance features of selected image pairs. The ‘language’ refers to the sentence that describes their similarity in terms of true gaze directions. Therefore, by semantically identifying the gaze difference of sample pairs with the help of the Differential Contrastive Training, we can fully leverage the semantic representation ability of CLIP to facilitate the primary gaze estimation model to characterize semantic gaze-related information.
@@
$\bullet$ We develop the potentials of CLIP in boosting gaze estimation performance with a novel Differential Contrastive Training strategy. Innovatively, we propose to establish the connections between languages and gaze differences of image pairs, which could provide new ideas for analogous quantified/regression tasks.
@@
Figure 2. Framework of our proposed DCGaze. It consists of two branch: Visual Appearance-aware branch and Semantic Difference-aware branch. In Visual Appearance-aware branch, the AFU is designed to adaptively enhance the gaze-related contents of primary appearance features with the guidance of the prior appearance features of CLIP’s image encoder. The DGR maps those enhanced features to final gazes via two regression heads. The Semantic Difference-aware branch selects several image pairs from sample batch and gives each of them a textual prompt that describes their gaze differences. Then, the prompts are fed into CLIP’s text encoder to capture text embeddings, which are then aligned with enhanced appearance features obtained by our Visual Appearance-aware branch.
2. Related Work
@@
Gaze estimation. With the recent developments in deep learning, appearance-based gaze estimation methods have become the mainstream. Researchers explored various CNN-based architectures (Zhang et al., 2017; Cheng et al., 2020a; Chen and Shi, 2019) to build the mapping between facial images and gaze directions. Zhang et al. (Zhang et al., 2015) firstly proposed a CNN-based gaze estimation network and the well-known MPIIFaceGaze dataset. With the introduction of Transformer (Vaswani et al., 2017), the Vision Transformer-based structures have started to be applied to gaze estimation and achieved promising performance (e.g., GazeTR (Cheng and Lu, 2022), oh et al.  (Oh et al., 2022), SUGE (Wang and Huang, 2024)). In the recent years, researchers dedicated themselves to inventing generalized gaze estimation methods, which would show robust performance on unseen users. Several methods leveraged a gaze redirection strategy to extend the datasets for generalized gaze estimation (Ganin et al., 2016; Xu and Lu, 2024). Cheng et al. (Cheng et al., 2022) introduced a method of purifying gaze features to improve the network’s generalization. Besides, some methods based on contrastive learning (Wang et al., 2022a)and uncertainty learning (Zhong et al., 2024; Wang and Huang, 2024) also demonstrated remarkable generalizability. In this paper, we leverage CLIP to improve discriminability and generalization of the gaze estimation network.
@@
Pre-trained Vision-language Models. Recently, the CLIP (Radford et al., 2021) trained on large-scale image-text pairs have attracted increasing attentions. Because of its powerful visual and semantic representation capabilities, CLIP has been transferred to various vision tasks, such as objection detection (Vidit et al., 2023), semantic segmentation (Zhou et al., 2023), and others (Li et al., 2022; Guo et al., 2024; Wu et al., 2024). Especially, CLIP also shows surprising capacities on quantified vision tasks, e.g., depth estimation (Zhang et al., 2022; Zeng et al., 2024), 3D hand poses estimation (Guo et al., 2023), etc. Moreover, researchers were attempting to take advantages of the pre-trained CLIP to predict gaze directions. Wang et al. (Wang et al., 2023) designed a linguistic description generator to produce text signals with coarse gaze directional cues. And then a cross-attention condenser was designed to finely recalibrate the visual and text representations, enhancing the learning quality of gaze features. Yin et al. (Yin et al., 2024) designed a feature separation loss by employing CLIP text encoder to generate gaze distractors from diverse language descriptions, which aims at purifying the gaze-relevant feature via pushing away it from gaze-irrelevant features. In this paper, we adapt CLIP to gaze estimation with an innovative collaborative enhancing strategy, in which the CLIP is regarded as an assistance to enhance the obtained gaze features.
3. Methodology
@@
To fully activate the potentials of CLIP to perceive gaze directions, we propose a novel Differential Contrastive Training (DCTrain) strategy, in which the CLIP is leveraged for unveiling gaze semantic differences of image pairs. Accordingly, we propose a Differential Contrastive Gaze Estimation network (DCGaze) as shown in Figure 2 which consists of a Visual Appearance-aware branch and a Semantic Difference-aware branch. The former is essentially a primary gaze estimation network, while the latter is to encourage the Visual Appearance-aware branch to extract gaze-related features with the help of image-language contrastive learning.
3.1. Visual Appearance-aware Branch
@@
The Visual Appearance-aware branch is composed of a Basic Appearance Feature Extractor and two innovative modules, a Adaptive Feature-refinement Unit (AFU) and a Double-head Gaze Regressor (DGR). The basic feature extractor captures the appearance feature of each facial image. The two proposed modules are designed to help extract informative and gaze-related appearance features.
3.1.1. Basic Appearance Feature Extractor
@@
We employ the remarkable CNN-Transformer architecture (Cheng and Lu, 2022) as the basic structure, where a ResNet is firstly adopted to acquire feature maps $f_{i}^{maps}\in\mathbb{R}^{W\times H\times C}$ of a given image $x_{i}\in X$ where $X=\{x_{1},x_{2},\dots,x_{n}\}$. Then those feature maps are reshaped into $W\times H$ patches $f_{i}^{p}\in\mathbb{R}^{\left(W\times H\right)\times C}$,which are treated as a series of $C$-dimensional visual tokens. After adding an extra learnable token $f_{i}^{token}\in\mathbb{R}^{1\times C}$, which is used to aggregate the features of all the patches, we feed them into a Transformer with a learnable position embedding $f_{i}^{pos}\in\mathbb{R}^{\left(1+W\times H\right)\times C}$. Overall, we get the final primary appearance feature $f_{i}^{pry}\in\mathbb{R}^{1\times C}$ as Eq. (1).
(1) $f_{i}^{pry}=\text{Transformer}\left(\left[{f_{i}^{token}};{f_{i}^{p}}\right]+{f_{i}^{pos}}\right)\left[0,:\right],$
@@
where [0,:][0,:][ 0 , : ] represents that the first row of the feature maps serves as the primary feature and [;][;][ ; ] denotes the concatenation operation.
3.1.2. Adaptive Feature-refinement Unit